You are standing 8 feet from a circular swimming pool. The distance to a
point of tangency is 16 feet. What is the radius of the pool?

Answer by fractalier(6550)

Make yourself a diagram such that the distance from the center of the pool to you is 8 + r and the distance to the tangent is 16. It should be a right triangle, since the radius meets a tangent at a right angle. Via the Pythagorean Theorem, we have
r^2 + 16^2 = (r + 8)^2
r^2 + 256 = r^2 + 16r + 64
256 = 16r + 64
16r = 192
r = 12 feet